# Navigating the Latent Space Dynamics of Neural Models

### v2

@@ 5 Related Work
Memorization and generalization in NNs NNs exhibit a rich spectrum of behaviors between memorization and generalization, depending on model capacity, regularization, and data availability [3, 57, 40]. In the case of extreme overparametrization, namely networks trained on few data points, it has been showed experimentally in [42, 56] and theoretically for sigmoid shallow AEs in [22] that AEs can memorize examples and implement associative memory mechanisms. [2, 51], A similar phenomenon has been observed in diffusion models in [48, 8] and analyzed in [24] Similarly, non gradient-based approaches such as Hopfield networks and their modern variants [19, 43] extend classical attractor dynamics to neural system that interpolate between memory-based and generalizing regimes. In our work, we show that AEs fall in general in the spectrum between memorization and generalization, depending on inductive biases that enforce contraction.
Contractive neural models: Different approaches have been proposed to regularize NNs in order to make them smoother and more robust to input perturbations and less prone to overfitting. Many of these regularization techniques either implicitly or explicitly promote contractive mappings in AEs: for example sparse AEs [37, 12], their denoising variant [52], and contractive AEs [44, 2] losses enforce learned maps to be contractive through the loss. Regularization strategies such as weight decay [28] favor as well contractive solutions for neural models, and they are incorporated in standardized optimizers such as AdamW [32]. All these approaches enforce either directly or indirectly the existence of fixed points and attractors in the proposed latent vector field representation.
Nonlinear operators spectral analysis In the context of image processing and 3D graphics, previous work has inspected generalization of spectral decompositions to nonlinear operators [5, 13, 10], focusing on one homogeneous operator (e.g., p-Laplacians). In our method, the fixed points computation of the vector field can be interpreted as the decomposition of the NN into a dictionary of signals
@@ 6 Conclusions and Discussion
In this manuscript we proposed to represent neural AEs as vectors fields, implicitly defined by iterating the decoding encoding map in the latent space. We showed that attractors in the latent vectors fields exists in practice due to inductive biases in the training regime which enforce local contractive solutions and they retain important properties of the model and the data, linking to memorization and generalization properties of the model. We show that the information stored in the weights can be retrieved without using any input data in vision foundation models and paths in the vector field are informative to characterize the learn distribution and detect distribution shifts.
Limitations and future works
Generalizing to arbitrary models Eq. 3 cannot be directly generalized to model trained with discriminative objectives such as a deep classifiers, or self supervised models, as the network is not invertible. However we note that point space the vector field is still defined in the output space, and can be simply by computing the residual $F(\mathbf{x})-y$. and in the neighborhood of an attractor, the relation in proposition 1 can still hold for different objectives.An alternative intriguing idea is the one to train a surrogate AE model in the latent space of the model of interest, which would be agnostic from the pretraining objective. Sparse AEs for mechanistic interpretability of large language models (LLMs) [12] fit in this category. Analyzing the associated latent vector field can shed light on features learned by SAEs and biases stored in their weights.
Learning dynamics: characterizing how attractors forms during training, under which conditions noise attractors converge to the training attractors, holds promise to use the proposed representation to study the learning dynamics of neural models to inspect finetuning of AE modules, such as low-rank adapters [20] and double descent [36].
Alignment of latent vector field: Finally, following recent findings in representation alignment [35, 11, 21] inspecting how latent vector fields of networks trained are related is an open question for future works, to possibly use this representations to compare different neural models.
Acknowledgments and Disclosure of Funding
This research was funded in whole or in part by the Austrian Science Fund (FWF) 10.55776/COE12. For open access purposes, the author has applied a CC BY public copyright license to any author accepted manuscript version arising from this submission. MF is supported as well by the MSCA IST-Bridge fellowship which has received funding from the European Union’s Horizon 2020 research and innovation program under the Marie Skłodowska-Curie grant agreement No 101034413. M.F. thanks C.Domine, V.Maiorca, I.Cannistraci, R.Cadei, B.Demirel, S.Vedula for insightful discussions.
@@ Appendix A Proofs and derivation
A.1 Theorem 3.1
We report here a detailed formulation of Theorem 3.1 alongside its proof.
@@
[…]
@@
=$\mathcal{M}=\left\{\mathbf{z}\in\mathbb{R}^{d}\;:\;f(\mathbf{z})=\mathbf{z}\right\}.$
@@
There exists an open, convex set $\Omega\subseteq\mathbb{R}^{d}$, with $\mathcal{M}\subseteq\Omega$, and a constant $10<L<$ 1, such that:
$\sup_{\mathbf{z}\in\Omega}\left\|J_{f}(\mathbf{z})\right\|_{\mathrm{op}}\leq L,$
where $J_{f}(\mathbf{z})\in\mathbb{R}^{d\times d}$ denotes the Jacobian of $f$ at $\mathbf{z}$, and $\|\cdot\|_{\mathrm{op}}$ denotes the operator norm (i.e., largest singular value). Therefore, $f$ is a contraction mapping on $\Omega$.
@@
Assume the encoder is deterministic, i.e., $q_{\phi}(\mathbf{z}|\mathbf{x})=\delta(\mathbf{z}-E(\mathbf{x}))$, so the latent marginal satisfies:
[…]
@@
[…]
@@ A.2 Proposition 1
$[(J_{f}(\mathbf{z})-\mathrm{Id})^{\top}+\alpha\mathrm{Id}]\mathbf{v}(\mathbf{z})=0$
@@
A.3 Proposition 3.2
We report here a detailed formulation of Theorem 3.2 alongside its proof.
@@ B.1 Bias at initialization
Figure 6: Contraction at initialization Variance preserving ratio at initialization of torchvision models: all models considered have a ratio $<1$, indicating that the map at initialization is contractive.
@@ B.4 Analysis speed of convergence
In figure 11 we report an empirical validation of proposition 3.2, for a 2 dimensional bottleneck convolution autoencoder trained on MNIST, by sampling points in the training set and measuring their convergence speed in terms of number of iteration. We remark that assumptions for the proposition don’t hold for this model, as the farer an initial condition is from an attractor the more higher order terms will dominate in the dynamics, making the evaluation of the spectrum of the Jacobian at the attractor insufficient. We observe nevertheless, that the convergence speed bounds still correlates, providing a looser estimate for the number of iteration needed to converge towards an attractors with error $\epsilon$.
@@ B.6 Memorization overfitting regime
In this section, we test the memorization capabilities of models in different (strong) overparametrization regimes, similarly as tested in [24] for diffusion models and in [42, 56] for the extremely overparametrized case (network trained on few samples). We remark that this case corresponds to an overfitting regime of the network, as opposed as underfitting (over-regularized) regime showed in the main paper. We train a convolutional autoencoder on subsamples of the CIFAR,MNIST,FashionMNIST datasets, with bottleneck dimension $k=128$ and weight decay $41e-$ 4. In Figure 13 we plot the memorization coefficient as a function of the dataset size, showing that networks trained on less data( strong overparametrization) tend to memorize more the training data.
@@ B.7 Data-free weight probing: additional results
We replicate the experiment in 4.1, on the larger XL version of the Stable diffusion architecture [46], observing that attractors from noise still form an informative dictionary of signals which scales better w.r.t, to a random orthogonal basis.
@@
Conv2D $3\to 128$, $3\times 3$, stride 1 $128\times 256\times 256$
@@ Appendix C Regularized autoencoder enforce contractiveness
In Table C, we list many AE variants including denoising AEs (DAEs) [52], sparse AEs (SAEs) [37], variational AEs (VAEs)  [26] and other variants [44, 2, 12] and show how their objectives enforce local contractive solutions around training points.
@@
R-Contractive AE [1] $\|J_{f_{\theta}}(\mathbf{x})\|_{F}^{2}$ Penalizes Jacobian of the full map $f_{\theta}=D_{\theta}\circ E_{\theta}$ Encourages stability of reconstructions: $f_{\theta}(\mathbf{x})\approx f_{\theta}(\mathbf{x}+\delta)$
Contractive AE [44] $\|J_{E_{\theta}}(\mathbf{x})\|_{F}^{2}=\sum_{i=1}^{k}\|J_{E_{i}}(\mathbf{x})\|^{2}$ Penalizes encoder Jacobian norm Explicitly enforces local flatness of encoder
Sparse AE (KL, sigmoid) [37] $\sum_{i=1}^{k}\mathrm{KL}(\rho\,\|\,\hat{\rho}_{i}),\quad\hat{\rho}_{i}=\mathbb{E}_{\mathbf{x}}[E_{i}(\mathbf{x})]$ Enforces low average activation under sigmoid Saturated units $\Rightarrow J_{E_{i}}(\mathbf{x})\approx 0\Rightarrow\|J_{E_{\theta}}(\mathbf{x})\|$ small
Sparse AE (L1, ReLU) [37] $\|E_{\theta}(\mathbf{x})\|_{1}$ Promotes sparsity of ReLU activations Inactive ReLU units have zero derivatives $\Rightarrow$ sparse Jacobian $J_{E_{\theta}}(\mathbf{x})$
Denoising AE [52, 2] $\mathbb{E}_{\tilde{\mathbf{x}}\sim q(\tilde{\mathbf{x}}|\mathbf{x})}\|\mathbf{x}-f_{\theta}(\tilde{\mathbf{x}})\|^{2}$ Reconstructs from noisy input For small noise: $f_{\theta}(\mathbf{x})-\mathbf{x}\approx\sigma^{2}\nabla\log p(\mathbf{x})$, a contractive vector field $J_{f_{\theta}}(\mathbf{x})$
Variational AE (VAE)[26] $\mathrm{KL}(q_{\theta}(\mathbf{z}|\mathbf{x})\,\|\,p(\mathbf{z}))$, where $q(\mathbf{z}|\mathbf{x})=\mathcal{N}(\mu(\mathbf{x}),\Sigma(\mathbf{x}))$ Encourages latent distribution to match prior Smooths $\mu(\mathbf{x}),\Sigma(\mathbf{x})$; penalizes sharp variations in encoder $J_{\mu}(\mathbf{x}),J_{\Sigma}(\mathbf{x})$
Masked AE (MAE)[15] $\mathbb{E}_{M}\left[\|M\odot(\mathbf{x}-f_{\theta}(M\odot\mathbf{x}))\|^{2}\right]$ Learns to reconstruct from partial input Promotes invariance to missing entries $\Rightarrow\|J_{f_{\theta}}(\mathbf{x})\|$ low

### v3

@@ 5 Related Work
Memorization and generalization in neural networks (NNs). NNs exhibit a rich spectrum of behaviors between memorization and generalization, depending on model capacity, regularization, and data availability (Arpit et al., 2017; Zhang et al., 2021; Power et al., 2022). In the case of extreme overparametrization, namely networks trained on few data points, it has been shown experimentally in (Radhakrishnan et al., 2020; Zhang et al., 2019) and theoretically for sigmoid shallow AEs in (Jiang and Pehlevan, 2020) that AEs can memorize examples and implement associative memory mechanisms. (Alain and Bengio, 2014; Vincent, 2011), A similar phenomenon has been observed in diffusion models in (Somepalli et al., 2023; Dar et al., 2023) and analyzed in (Kadkhodaie et al., 2023) Similarly, non gradient-based approaches such as Hopfield networks and their modern variants (Hopfield, 1982; Ramsauer et al., 2020) extend classical attractor dynamics to neural systems that interpolate between memory-based and generalizing regimes. In our work, we show that AEs fall in general in the spectrum between memorization and generalization, depending on inductive biases that enforce contraction.
Contractive neural models. Different approaches have been proposed to regularize NNs in order to make them smoother and more robust to input perturbations and less prone to overfitting. Many of these regularization techniques either implicitly or explicitly promote contractive mappings in AEs: for example, sparse AEs (Ng et al., 2011; Gao et al., 2024), their denoising variant (Vincent et al., 2008), and contractive AEs (Rifai et al., 2011; Alain and Bengio, 2014) losses enforce learned maps to be contractive through the loss. Regularization strategies such as weight decay (Krogh and Hertz, 1991) favor as well contractive solutions for neural models, and they are incorporated in standardized optimizers such as AdamW (Loshchilov, 2017). All these approaches enforce either directly or indirectly the existence of fixed points and attractors in the proposed latent vector field representation.
Neural networks as dynamical systems. Distinct lines of work have interpreted neural networks as dynamical systems. Neural ODEs (Chen et al., 2018) view depth as continuous time and model hidden states via differential equations, while deep equilibrium models (Bai et al., 2019) characterize predictions as fixed points of implicit dynamics. Closer to our work, Radhakrishnan et al. (2020) interpret overparameterized autoencoders as dynamical systems acting in the input space, implementing associative memory. In contrast, we show that any autoencoder induces a latent vector field, and we link its properties to generalization, memorization, and the characterization of the learned distribution.
Nonlinear operators spectral analysis. In the context of image processing and 3D graphics, previous work has inspected generalization of spectral decompositions to nonlinear operators (Bungert et al., 2021; Gilboa et al., 2016; Fumero et al., 2020), focusing on one homogeneous operator. In our work, fixed points of Eq. 3 can be interpreted as the decomposition of the NN into a dictionary of signals.
@@ 6 Conclusions and Discussion
In this work we proposed to represent neural AEs as vectors fields, implicitly defined by iterating the autoencoding map in the latent space. We showed that (i) attractors in the latent vectors field exists in practice due to inductive biases in the training regime which enforce local contractions; (ii) they retain key properties of the model and the data, linking to memorization and generalization regimes of the model; (iii) knowledge stored in the weights can be retrieved without access to input data in vision foundation models; (iv) paths in the vector field inform on the learned distribution and its shifts.
Limitations and future works. Generalizing to arbitrary models: Eq. 3 cannot be directly generalized to model trained with discriminative objectives such as a deep classifiers, or self supervised models, as the network is not invertible. However we note that point space the vector field is still defined in the output space, and can be simply by computing the residual $F(\mathbf{x})-y$ and in the neighborhood of an attractor, the relation in proposition 3.1 can still hold for different objectives.An alternative intriguing idea is the one to train a surrogate AE model in the latent space of the model of interest, which would be agnostic from the pretraining objective. Sparse AEs for mechanistic interpretability of large language models (LLMs) (Gao et al., 2024) fit in this category. Analyzing the associated latent vector field can shed light on features learned by SAEs and biases stored in their weights. Learning dynamics: characterizing how attractors forms during training, under which conditions noise attractors converge to the training attractors, holds promise to use the proposed representation to study the learning dynamics of neural models to inspect finetuning of AE modules, such as low-rank adapters (Hu et al., 2022) and double descent (Nakkiran et al., 2021). Alignment of latent vector fields: finally, following recent findings in representation alignment (Moschella et al., 2022; Fumero et al., 2024; Huh et al., 2024) inspecting how latent vector fields of networks trained are related is an open question for future works, to possibly use this representations to compare different neural models.
@@ Appendix A Proofs and derivation
A.1 Theorem 1
@@
$q(\mathbf{z})=\int p_{\mathrm{data}}(\mathbf{x})\,q_{\phi}(\mathbf{z}\mid\mathbf{x})\,d\mathbf{x}.$
@@
$\mathcal{M}=\left\{\mathbf{z}\in\mathbb{R}^{d}\;:\;f(\mathbf{z})=\mathbf{z}\right\}.$
@@
There exists an open, convex set $\Omega\subseteq\mathbb{R}^{d}$, with $\mathcal{M}\subseteq\Omega$, and a constant $0<L<1$, such that:
$\sup_{\mathbf{z}\in\Omega}\left\|J_{f}(\mathbf{z})\right\|_{\sigma}\leq L,$
where $J_{f}(\mathbf{z})\in\mathbb{R}^{d\times d}$ denotes the Jacobian of $f$ at $\mathbf{z}$, and $\|\cdot\|_{\sigma}$ denotes the spectral norm (i.e., largest singular value). Therefore, $f$ is a contraction mapping on $\Omega$.
@@
For example, assume the encoder is deterministic, i.e., $q_{\phi}(\mathbf{z}|\mathbf{x})=\delta(\mathbf{z}-E(\mathbf{x}))$, so the latent marginal satisfies:
$q(\mathbf{z})=\int p_{\mathrm{data}}(\mathbf{x})\,\delta(\mathbf{z}-E(\mathbf{x}))\,d\mathbf{x}.$
@@
Step 1: Contraction and convergence. By Assumption 3, $f:\Omega\to\Omega$ is a contraction mapping with contraction constant $L<1$. Since $\Omega$ is convex and hence complete under $\|\cdot\|$, Banach’s fixed-point theorem applies. Therefore:
@@
We report here a detailed formulation of Proposition 3.1 alongside its proof.
@@
$(J_{f}(\mathbf{z})-\mathrm{Id})^{\top}+\alpha\mathrm{Id}]\mathbf{v}(\mathbf{z})=0$
@@
A.3 Attractors connects to generalization error
We report here a detailed formulation of Proposition 3.2 alongside its proof.
Proposition A.8 (Attractors as prototypes for generalization).
Let $\mathbf{Z}^{\star}\subset\Omega\subset\mathbb{R}^{d}$ be a (finite) dictionary of attractors of $f=E\circ D$ in a neighborhood $\Omega$ of latent space, and let $\Pi:\Omega\to\mathbf{Z}^{\star}$ denote a (measurable) nearest-attractor map $\Pi(\mathbf{z})\in\arg\min_{\mathbf{u}\in\mathbf{Z}^{\star}}\|\mathbf{z}-\mathbf{u}\|_{2}$ (with any fixed tie-breaking rule). If the decoder $D$ is $L_{D}$-Lipschitz on $\Omega$, then for any test point $\mathbf{x}$ with $\mathbf{z}=E(\mathbf{x})\in\Omega$:
$\big\|\mathbf{x}-F(\mathbf{x})\big\|_{2}^{2}\;\leq\;\underbrace{\big\|\mathbf{x}-D(\Pi(\mathbf{z}))\big\|_{2}^{2}}_{\text{prototype error}}\;+\;\underbrace{L_{D}^{2}\,\big\|\mathbf{z}-\Pi(\mathbf{z})\big\|_{2}^{2}}_{\text{coverage error}}.$
Moreover, if $\mathbf{Z}^{\star}$ is an $\varepsilon$-cover of $\mathrm{supp}\,q$ (the latent marginal), then for all $\mathbf{x}$ with $E(\mathbf{x})\in\mathrm{supp}\,q$,
$\big\|\mathbf{x}-F(\mathbf{x})\big\|_{2}^{2}\;\leq\;\big\|\mathbf{x}-D(\Pi(E(\mathbf{x})))\big\|_{2}^{2}\;+\;L_{D}^{2}\varepsilon^{2},$
and the same bound holds in expectation over $\mathbf{x}\sim p_{\mathrm{test}}$ whenever $E(\mathbf{x})\in\mathrm{supp}\,q$ almost surely.
We make the following mild assumptions:
  • (A1)
    $\mathbf{Z}^{*}\subset\Omega$ is a (finite) set of attracting fixed points of $f=E\circ D$ contained in an open neighborhood $\Omega\subset\mathbb{R}^{d}$.
  • (A2)
    $D$ is $L_{D}$-Lipschitz on $\Omega$, i.e., $\|D(\mathbf{z})-D(\mathbf{u})\|\leq L_{D}\|\mathbf{z}-\mathbf{u}\|_{2}$ for all $\mathbf{z},\mathbf{u}\in\Omega$.
  • (A3)
    $\Pi:\Omega\to\mathbf{Z}^{\star}$ is any measurable selection of nearest points, e.g. $\Pi(\mathbf{z})\in\arg\min_{\mathbf{u}\in\mathbf{Z}^{\star}}\|\mathbf{z}-\mathbf{u}\|_{2}$ .
Proof of Proposition A.8.
Let $\mathbf{x}$ be any test point with $\mathbf{z}=E(\mathbf{x})\in\Omega$. Add and subtract $D(\Pi(\mathbf{z}))$:
$\mathbf{x}-D(E(\mathbf{x}))=\underbrace{\mathbf{x}-D(\Pi(\mathbf{z}))}_{\text{prototype error}}\;+\;\underbrace{D(\Pi(\mathbf{z}))-D(E(\mathbf{x}))}_{\text{decoder distortion}}.$
Taking norms and using Cauchy-Schwarz inequality gives:
$\|\mathbf{x}-D(E(\mathbf{x}))\|_{2}^{2}\;\leq\;\|\mathbf{x}-D(\Pi(\mathbf{z}))\|_{2}^{2}\;+\;\|D(\Pi(\mathbf{z}))-D(E(\mathbf{x}))\|_{2}^{2}.$
By (A2), $\|D(\Pi(\mathbf{z}))-D(E(\mathbf{x}))\|\leq L_{D}\|\Pi(\mathbf{z})-\mathbf{z}\|_{2}$, which yields:
$\|\mathbf{x}-F(\mathbf{x})\|_{2}^{2}\;\leq\;\|\mathbf{x}-D(\Pi(\mathbf{z}))\|_{2}^{2}+L_{D}^{2}\|\mathbf{z}-\Pi(\mathbf{z})\|_{2}^{2}.$
This proves the first inequality. If $\mathbf{Z}^{\star}$ is an $\varepsilon$-cover of $\mathrm{supp}\,q$, then for any $\mathbf{z}\in\mathrm{supp}\,q$ we have $\|\mathbf{z}-\Pi(\mathbf{z})\|\leq\varepsilon$, hence
$\|\mathbf{x}-F(\mathbf{x})\|_{2}^{2}\;\leq\;\|\mathbf{x}-D(\Pi(E(\mathbf{x})))\|_{2}^{2}+L_{D}^{2}\,\varepsilon^{2},$
for all $\mathbf{x}$ with $E(\mathbf{x})\in\mathrm{supp}\,q$. Taking expectations (when $E(\mathbf{x})\in\mathrm{supp}\,q$ almost surely) gives the stated bound. $\square$
Remarks.
(i) The bound decomposes test error into a prototype term, $\|\mathbf{x}-D(\Pi(E(\mathbf{x})))\|_{2}^{2}$, and a coverage term, $L_{D}^{2}\|\mathbf{z}-\Pi(\mathbf{z})\|_{2}^{2}$. (ii) Under the assumptions of Theorem 1, $\mathbf{Z}^{*}$ are attractors with basins; then $\Pi$ aligns with the destination of latent dynamics, tying the algebraic bound to the memorization–generalization picture.
A.4 Attractors in simple networks
In this section we characterize attractors in linear and homogeneous networks, showing that in this simple setting is possible to prove converge speed to attractors of the iteration in Eq 3. Linear maps. In the linear case, the encoding map $E$ and decoding $D$ are parametrized by matrices $\mathbf{W}_{1}\in\mathbb{R}^{N\times k}$ and $\mathbf{W}_{2}\in\mathbb{R}^{k\times N}$ and the only fixed point of the map corresponds to the origin if the network is bias free or to a shift of it. The rate of convergence of the iteration in Eq 3 is established by the spectrum of $\mathbf{W}_{2}^{T}\mathbf{W}_{1}$, and the iteration is equivalent to shrinking the input in the direction of the eigenvectors with associated eigenvalue $\lambda<1$. In case the eigenvectors of the trained AEs are aligned with the optimal solution given by the Principal Component Analysis (PCA) decomposition of the data $\mathbf{X}=\mathbf{\Phi}\mathbf{\Lambda}\mathbf{\Phi}^{*}$, then the latent vector field vanished as the mapping is isometric and no contraction occurs.
Homogeneous maps. In the case of homogeneous neural networks, the network satisfies $F(c\mathbf{x})=c^{\alpha}F(\mathbf{x})$ with $c\in\mathbb{R}$ for some $\alpha$. For example, this holds for ReLU networks without biases with $\alpha=1$, which learn a piecewise linear mapping. The input-output mapping can be rewritten as:
$F(\mathbf{x})=J_{F}(\mathbf{x})\mathbf{x}$ (6)
A similar observation was made in for denoising networks in [Kadkhodaie et al., 2023], we remark here its generality.
This equality implies that we can rewrite Equation 3 as :
Proposition A.9 (informal, proof in Appendix A.5).
The error $e_{t}=\|\mathbf{z}_{t}-\mathbf{z}^{*}\|$ converge exponentially to an $\epsilon$ depending on the spectral norm $\|J_{f}(\mathbf{z}^{*})\|_{\sigma}$, according to the formula $\frac{\log(\frac{\epsilon}{\|\mathbf{e}_{0}\|})}{\log(\|J_{f}(\mathbf{z}^{*})\|_{\sigma})}$ that provides an estimate for the number of iterations $T$ to converge to the attractor.
In Figure 11 in the Appendix, we measure how well the convergence formula predicts the measured number of iterations to converge to an attractor, on a fully non linear AE with biases, where the assumptions of the current section are less likely to hold. The error in the estimate will be higher, when the initial condition $\mathbf{z}_{0}$ is far from the attractor, as higher order terms dominate the dynamics, and the first order Taylor approximation accumulates more error.
A.5 Proof of Proposition A.9
We report here a detailed formulation of Proposition A.9 alongside its proof.
@@ B.1 Bias at initialization
Figure 6: Contraction at initialization Variance preserving ratio at initialization of torchvision models: all models considered have a ratio <1<1, indicating that the map at initialization is contractive.
@@ B.4 Analysis speed of convergence
In figure 11 we report an empirical validation of proposition A.9, for a 2 dimensional bottleneck convolution autoencoder trained on MNIST, by sampling points in the training set and measuring their convergence speed in terms of number of iteration. We remark that assumptions for the proposition don’t hold for this model, as the farer an initial condition is from an attractor the more higher order terms will dominate in the dynamics, making the evaluation of the spectrum of the Jacobian at the attractor insufficient. We observe nevertheless, that the convergence speed bounds still correlates, providing a looser estimate for the number of iteration needed to converge towards an attractors with error $\epsilon$.
@@ B.6 Memorization overfitting regime
In this section, we test the memorization capabilities of models in different (strong) overparametrization regimes, similarly as tested in [Kadkhodaie et al., 2023] for diffusion models and in [Radhakrishnan et al., 2020, Zhang et al., 2019] for the extremely overparametrized case (network trained on few samples). We remark that this case corresponds to an overfitting regime of the network, as opposed as underfitting (over-regularized) regime showed in the main paper. We train a convolutional autoencoder on subsamples of the CIFAR,MNIST,FashionMNIST datasets, with bottleneck dimension $k=128$ and weight decay 1e41e-4. In Figure 13 we plot the memorization coefficient as a function of the dataset size, showing that networks trained on less data( strong overparametrization) tend to memorize more the training data.
@@ B.7 Data-free weight probing: additional results
We replicate the experiment in 4.1, on the larger XL version of the Stable diffusion architecture [Rombach et al., 2022], observing that attractors from noise still form an informative dictionary of signals which scales better w.r.t, to a random orthogonal basis.
@@
Conv2D $\to 128, 3×3$$\times 3$, stride 1 $128\times 256\times 256$
@@ Appendix C Regularized autoencoder enforce contractiveness
In Table C, we list many AE variants including denoising AEs (DAEs) [Vincent et al., 2008], sparse AEs (SAEs) [Ng et al., 2011], variational AEs (VAEs)  [Kingma et al., 2013] and other variants [Rifai et al., 2011, Alain and Bengio, 2014, Gao et al., 2024] and show how their objectives enforce local contractive solutions around training points.
@@
R-Contractive AE [Alain and Olivier, 2013] $\|J_{f_{\theta}}(\mathbf{x})\|_{F}^{2}$ Penalizes Jacobian of the full map $f_{\theta}=D_{\theta}\circ E_{\theta}$ Encourages stability of reconstructions: $f_{\theta}(\mathbf{x})\approx f_{\theta}(\mathbf{x}+\delta)$
Contractive AE [Rifai et al., 2011] $\|J_{E_{\theta}}(\mathbf{x})\|_{F}^{2}=\sum_{i=1}^{k}\|J_{E_{i}}(\mathbf{x})\|^{2}$ Penalizes encoder Jacobian norm Explicitly enforces local flatness of encoder
Sparse AE (KL, sigmoid) [Ng et al., 2011] $\sum_{i=1}^{k}\mathrm{KL}(\rho\,\|\,\hat{\rho}_{i}),\quad\hat{\rho}_{i}=\mathbb{E}_{\mathbf{x}}[E_{i}(\mathbf{x})]$ Enforces low average activation under sigmoid Saturated units $\Rightarrow J_{E_{i}}(\mathbf{x})\approx 0\Rightarrow\|J_{E_{\theta}}(\mathbf{x})\|$ small
Sparse AE (L1, ReLU) [Ng et al., 2011] $\|E_{\theta}(\mathbf{x})\|_{1}$ Promotes sparsity of ReLU activations Inactive ReLU units have zero derivatives $\Rightarrow$ sparse Jacobian $J_{E_{\theta}}(\mathbf{x})$
Denoising AE [Vincent et al., 2008, Alain and Bengio, 2014] $\mathbb{E}_{\tilde{\mathbf{x}}\sim q(\tilde{\mathbf{x}}|\mathbf{x})}\|\mathbf{x}-f_{\theta}(\tilde{\mathbf{x}})\|^{2}$ Reconstructs from noisy input For small noise: $f_{\theta}(\mathbf{x})-\mathbf{x}\approx\sigma^{2}\nabla\log p(\mathbf{x})$, a contractive vector field $J_{f_{\theta}}(\mathbf{x})$
Variational AE (VAE)[Kingma et al., 2013] $\mathrm{KL}(q_{\theta}(\mathbf{z}|\mathbf{x})\,\|\,p(\mathbf{z}))$, where $q(\mathbf{z}|\mathbf{x})=\mathcal{N}(\mu(\mathbf{x}),\Sigma(\mathbf{x}))$ Encourages latent distribution to match prior Smooths $\mu(\mathbf{x}),\Sigma(\mathbf{x})$; penalizes sharp variations in encoder $J_{\mu}(\mathbf{x}),J_{\Sigma}(\mathbf{x})$
Masked AE (MAE)[He et al., 2022] $\mathbb{E}_{M}\left[\|M\odot(\mathbf{x}-f_{\theta}(M\odot\mathbf{x}))\|^{2}\right]$ Learns to reconstruct from partial input Promotes invariance to missing entries $\Rightarrow\|J_{f_{\theta}}(\mathbf{x})\|$ low